What type of questions indicate the interpretation of multiplication as scaling?
1
Expert's answer
2020-06-18T18:13:51-0400

If it is related to the multiplication of any number less than 1, then the answer becomes smaller, e.g. 5 x 1/2 =5/2 =2.5.

If it is multiplied by the number greater than 1, the answer becomes greater, e.g. 5 x 2 =10.